How many times more powerful was the earthquake in Alaska, with a magnitude of 9.2 on the Richter scale, compared to the San Francisco earthquake, which had a magnitude of 7.9 on the Richter scale?
Answer by math_tutor2020(3838)


Let's do an example problem before tackling this current one.

Let's say we want to compare a 5.0 to a 6.0 on the Richter scale.
The second value is 10 times greater because 6-5 = 1 and 10^1 = 10.

Another example:
Comparing a 4.0 to a 7.0 means 7-4 = 3 and 10^3 = 1000. So the 7.0 magnitude is 1000 times greater than the 4.0


Now onto the current problem:
As the examples above show, we first subtract the magnitudes to get 9.2 - 7.9 = 1.3

Then use a calculator to get
10^(1.3) = 19.9526
which is approximate
The decimal digits go on forever without a known pattern. This decimal value is irrational.


Answer: 19.9526 roughly
Round this however you need